OpenStudy (anonymous):

For which values of n is 1n+2n+3n even

OpenStudy (freckles):

well 1n+2n+3n=what?

OpenStudy (freckles):

1+2+3=?

OpenStudy (zzr0ck3r):

if you put in an odd number, then 1*odd is odd, 2*odd is even and 3*odd is odd so we get odd+even+odd = even if you put in even n then 1*even = even, 2*even = even and 3* even = even so we get even + even + even = even so the equation is true for all n

OpenStudy (zzr0ck3r):

that should say so the equation is even for all n*

OpenStudy (zzr0ck3r):

just in case you wanted another way to think about it:)

OpenStudy (unklerhaukus):

\[ \{n|1n+2n+3n=2m|m\in\mathbb N\}\\ \{n|6n=2m|m\in\mathbb N\}\\ \{n|3n=m|m\in\mathbb N\}\\ \{n|n=m/3|m\in\mathbb N\}\\ \]